OpenStudy (anonymous):

Which of the following points are NOT on the graph of y = -2x + 3?

OpenStudy (anonymous):

What are the options?

OpenStudy (anonymous):

1,1 -1,5 -2,3 2,-1

OpenStudy (anonymous):

It's -2,3. It took me a while to read the graph coordinates, but there's no -2,3

OpenStudy (anonymous):

thanks zeinas !

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

Which of the following points are NOT on the graph of y = -4x + 4? 1,0 2,4 0,4 -1,8

OpenStudy (anonymous):

(2,4). It's completely off the graph. And one way that you can do this is to put the points in the equation, and the one that doesn't make it equal to each other is wrong.

OpenStudy (anonymous):

Which of the following points are NOT on the graph of y = x + 2? -1,1 1,3 3,5 5,9 My answer was 5,9 ?

OpenStudy (anonymous):

That's right :D

OpenStudy (anonymous):

this one confused me a little bit my answer was 9,7 The answers are 9,7 -3,-1 -6,-3 6,3

OpenStudy (anonymous):

What's the equation for the graph?

OpenStudy (anonymous):

\[y=\frac{ 2 }{ 3 }x+1\]

OpenStudy (anonymous):

Well it could not be 9,7 because 7=2/3(9)+1 7=6+1 7=7 But when you try the last point (6,3) we get 3=2/3(6)+1 3=4+1 3=5 and that is wrong so its d.

OpenStudy (anonymous):

thanks (: what about the y=x^2-1 0, -1 1, 0 -1, 0 0, 0 is it C ?

OpenStudy (anonymous):

0=(-1)^2-1 0=1-1 0=0 so it fits But D is 0=(0)^2-1 0=-1 It's D because it doesn't fit.

OpenStudy (anonymous):

thanks.Which of the following points are NOT on the graph of y = 2x? -2, 0 2, 4 4, 8 0, 0

OpenStudy (anonymous):

It's a. 0=2(-2) 0=-4

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

No problem. :)